Please answer the following question correctly and completely. Please write neatly and show all work.

2. In a study of author productivity (Lotkas Test, Collection Mgmt., 1982: 111-118), a large number of authors were classi
Answer #1
# of papers Frequecy Rel . Freq
1 784 0.59893048
2 204 0.15584416
3 127 0.09702063
4 50 0.0381971
5 33 0.02521008
6 28 0.02139037
7 19 0.0145149
8 19 0.0145149
9 6 0.00458365
10 7 0.00534759
11 6 0.00458365
12 7 0.00534759
13 4 0.00305577
14 4 0.00305577
15 5 0.00381971
16 3 0.00229183
17 3 0.00229183
sum 1309 1

(a) Plat least spapers) = P(x>5) = 1 - P(x45) - 1-T0.5989 +0.1558 + 0.097 0 + 0.0382 = 10.1101 (6) P(x <S) = 0.5989 + 0.1558